OpenStudy (anonymous):

when finding a perimeter of a rectangle. if one side is the square root of 27 and the other is square root of 108. what would the exact perimeter be?

OpenStudy (anonymous):

Sqrt(27) + sqrt(108) + sqrt(27) + sqrt(108) Then you need to simplify each radical Do you know who to do that? sqrt(27) = 2sqrt(3) What is sqrt(108) =

OpenStudy (anonymous):

i show its a decimal is that right?

OpenStudy (anonymous):

No.. you can't do it that way. If they want "exact" Have you been working on radicals and simplifying them?

OpenStudy (anonymous):

Do you know how to do a factor tree?

OpenStudy (anonymous):

yeah we have to find the lcd or gcf

OpenStudy (anonymous):

27 = 3 * 3 * 3 108 = 2 * 54 = 2 * 2 * 27 = 2 * 2 * 3 * 9 = 2 * 2 * 3 * 3 * 3 So 108= 2 * 2 * 3 * 3 *3 Doyou understand this?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

OK the way you simplify a radical is - circle groups of two that are alike. FOr example. 27 = 3 * 3 * 3 you would circle the first two 3's and they go outside the sign with the uncircled 3 left underneath. sqrt(27) = 3sqrt(3) 108 = 2 * 2 * 3 * 3 * 3 you would circle the 2's and the 3's and there would be one 3 left over. You multiply things that come out, and multiply things that are left underneath. sqrt(108) = 6sqrt(3) So 3sqrt(3) + 6sqrt(3) + 3 sqrt(3) + 6sqrt(3) = 18sqrt(3) That is your answer...

OpenStudy (anonymous):

wow i really like that break down thank you

OpenStudy (anonymous):

You are welcome